Ali walks 680m south 780m east 80m south find the distance and the displacement?
1
Expert's answer
2020-03-17T09:22:53-0400

The distance


"d=680+780+80=1540\:\rm m."

The displacement


"s=\sqrt{(680+80)^2+780^2}=1089\:\rm m."